Suppose a population is composed of only 3 measures: 1, 2, and 3. The possible samples of size 2 can be draw from this population. List all the possible sample size 2 when repetition is allowed or with replacement.

1
Expert's answer
2022-03-19T04:09:39-0400

Possible samples are

"\left( 1,1 \right) ,\left( 1,2 \right) ,\left( 1,3 \right) \\\left( 2,1 \right) ,\left( 2,2 \right) ,\left( 2,3 \right) \\\left( 3,1 \right) ,\left( 3,2 \right) ,\left( 3,3 \right)"